OpenStudy (anonymous):

solve the inequality -5x ≥ -10 or 2(x + 3) ≥ 14

OpenStudy (anonymous):

not true. 2(x+3)>=14 2x+6>=14 2x>=8 x>=4

OpenStudy (anonymous):

3x + 5y = 51 y=2x + 5